OpenStudy (anonymous):

Differentiate the function. F(x)=ln rad x. I switched it to ln x^(1/2) since I'm not good with radicals

ganeshie8 (ganeshie8):

thats a good start, heard of chain rule before ?

OpenStudy (anonymous):

yup is that what needs to be done for x^(1/2)? In my book it says the answer is F'(x)=1/(2x) but I have no idea how they got that

ganeshie8 (ganeshie8):

Yeah you can avoid chain rule for now by using log properties

ganeshie8 (ganeshie8):

\[\large \rm F(x) = \ln \sqrt{x} = \ln x^{1/2} = \dfrac{1}{2} \ln x\]

ganeshie8 (ganeshie8):

Now differentiate

OpenStudy (anonymous):

ok I got it thank you :)

ganeshie8 (ganeshie8):

yw:)